# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v2

@@ 5.1 Effectiveness of Shifts-aware Reward
Effectiveness in Complex Tasks. We also conducted experiments on complex tasks. Fig. 3 shows a visual comparison in the NeoRL task (details in Sec. 5.2.2). SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data, whereas the previous state-of-the-art method, MOBILE, imposes weaker constraints. Furthermore, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with the Offline Data and remains tightly concentrated, indicating the SAMBO’s effectiveness. In contrast, the Real Data from MOBILE deviates significantly from both the Offline and Model Data, still suffering from distribution shift, leading to diminished performance. We attribute this discrepancy to SAMBO’s robust handling of model bias and policy shift, ensuring that the Real Data, Model Data, and Offline Data remain closely aligned, effectively mitigating the distribution shift.
5.2 Benchmark Results
5.2.1 D4RL
Datasets. In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012). The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert). The “v2” versions of datasets are employed for assessment.
Baselines. We evaluate SAMBO against several baseline methods, including model-free methods such as BC (Bain & Sammut, 1995; Ross et al., 2011), which learns the behavior policy, CQL (Kumar et al., 2020), TD3+BC (Fujimoto & Gu, 2021), and EDAC (An et al., 2021), as well as model-based methods such as MOPO (Yu et al., 2020), COMBO (Yu et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023).
Table 2: Normalized average returns in NeoRL benchmark over 4 random seeds. The results for MOBILE are reported from the original paper, and MOBILE represents the results obtained by executing the official implementation.
Task Name BC CQL TD3+BC EDAC MOPO MOBILE MOBILE SAMBO (Ours)
HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.7 54.8 39.8 $\pm$ 2.6
Hopper-L 15.1 16.0 15.8 18.3 6.2 17.4 18.2 19.4 $\pm$ 1.3
Walker2d-L 28.5 44.7 43.0 40.2 11.6 37.6 23.7 46.1 $\pm$ 5.7
HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 77.8 79.0 62.5 $\pm$1.6
Hopper-M 51.3 64.5 70.3 44.9 1.0 51.1 43.4 67.6 $\pm$ 12.2
Walker2d-M 48.7 57.3 58.6 57.6 39.9 62.2 60.1 63.4 $\pm$ 4.0
HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 83.0 71.8 71.6 $\pm$ 13.9
Hopper-H 43.1 76.6 75.3 52.5 11.5 87.8 42.3 93.9 $\pm$ 15.2
Walker2d-H 72.6 75.3 69.6 75.5 18.0 74.9 71.9 79.2 $\pm$ 0.6
@@ 5.2.1 D4RL
Figure 4: Performance comparison of Hopper tasks in NeoRL benchmark.
Comparison Results. Results are presented in Table 1. Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al., 2020). SAMBO achieves strong performance on random datasets and demonstrates competitive performance across medium, medium-replay, and medium-expert datasets. This performance aligns with our theoretical analysis. By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy. This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy. In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy. Conversely, a conservative approach may impede the learning of optimal policy in such contexts. For high-quality datasets, the advantages of aggressive exploration tend to diminish. Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively. This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
5.2.2 NEORL
@@
Datasets. NeoRL (Qin et al., 2022) is a benchmark designed to emulate real-world scenarios by generating datasets through a more conservative policy, aligning closely with real-world data-collection scenarios. The datasets, characterized by their narrow scope and limited coverage, present a substantial challenge for offline RL algorithms. In this work, we examine nine datasets across three environments (HalfCheetah-v3, Hopper-v3, Walker2d-v3) and three dataset quality levels (L, M, H), corresponding to low, medium, and high-quality datasets. NeoRL offers different numbers of trajectories for each task (100, 1000, and 10000), and we consistently selected 1000 trajectories in our test.
@@
Baselines. We compare SAMBO with baseline methods that align with the d4rl experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time.
@@
Comparison Results. Results are presented in Table 2 and Figure 4. We compare the learning curves of model-based approaches in the NeoRL benchmark. Since MOPO does not have an official implementation on NeoRL, we report its score based on the original NeoRL paper. The experiments show that Our algorithm achieves superior or competitive performance on the more challenging and realistic NeoRL datasets, which can be seen as lower-quality datasets compared to the D4RL datasets. In particular, SAMBO demonstrates exceptional performance in the Walker2d and Hopper environments, which, in contrast to the HalfCheetah environment, exhibit higher control complexity, more intricate action dimensions, and greater training difficulty. These offline datasets more accurately capture the dynamic control challenges encountered in real-world scenarios, making them significantly demanding and highlighting that SAMBO has strong potential for real-world applications. More experiments and discussions are provided in Appendix D.3.
(a) Ablation in halfcheetah-random dataset
(b) Tuning for $\alpha$
(c) Tuning for $\beta$
(d) Ablation study in Walker2d-H dataset
Figure 5: Illustration of hyperparameter experiments in D4RL benchmark and NeoRL benchmark. (a) and (d) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward. (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness. (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
5.3 Ablation Studies
We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 5.
@@ 5.3 Ablation Studies
Ablation. Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Halfcheetah-random task in D4RL benchmark and Walker2d-H task in NeoRL benchmark. As shown in Figure. 5 (a) and (d), we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent.
Figure. 5 (a) show that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates. This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration. However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance. This negative effect is particularly evident in the narrow Walker2d-H dataset, as seen in Figure 5 (d). Conversely, the wo/ps ablation addresses model bias but neglects policy shift adjustments. While this reduces value misestimation, the algorithm remains constrained by the behavior policy due to the lack of policy shift modification. Consequently, wo/ps is less effective in discovering a more optimal policy. As shown in Figure 5 (a), the performance of wo/ps is inferior to sambo, and Figure 5 (d) illustrates that wo/ps converges slowly compared to sambo. In contrast, SAMBO incorporates not only model bias adjustment but also policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance.
Hyperparameter Tuning. SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$. The model bias adjustment coefficient $\alpha$ adjusts model-generated data, with larger values overly penalizing exploration and smaller values inducing over-optimism, both harming performance. To validate this, we fixed $\beta=0.01$ and varied $\alpha$ in the halfcheetah task. As shown in Fig. 5 (b), increasing $\alpha$ from 0 to 0.05 initially improves performance but later hinders it due to excessive conservativeness, highlighting the need for a balanced $\alpha$ in policy learning.
The policy shift coefficient $\beta$ regulates modifications to offline data, controlling deviation from the behavior policy. Larger $\beta$ values enhance exploration by encouraging greater deviation from the behavior policy, but also increase the risk of unsafe or suboptimal actions. As shown in Fig. 5 (c), the algorithm exhibits performance fluctuations as $\beta$ increases from 0 to 0.1. Low $\beta$ may lead to suboptimal solutions, while high $\beta$ can induce excessive aggressiveness and instability. Therefore, $\beta$ also needs to be set appropriately.
6 Conclusion and Discussion
@@
In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift. Model bias results in distorted objectives, while policy shift obstructs the convergence of policies. We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR) and apply it within the model-based offline setting, thereby proposing our SAMBO-RL algorithm. Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for classifier estimation. Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems. Our novel analysis and results related to policy shifts in this context open new avenues for better exploiting offline datasets and overcoming the implicit restrictions imposed by behavior policies.
Impact Statement
This study addresses the distribution shift problem in offline model-based reinforcement learning (RL), with the potential to facilitate the deployment of RL algorithms in real-world applications, particularly in scenarios where data collection is expensive. The research aims to enhance the robustness and reliability of RL systems in real world. However, the direct deployment of RL systems in high-risk domains, such as autonomous driving, may introduce substantial risks, especially if algorithms fail to adapt to unforeseen distribution shifts in real-world environments. Furthermore, the application of offline RL models, particularly in sensitive areas like healthcare, may raise significant ethical concerns. In such domains, the training data may not fully represent all patient groups. If distribution shift issues are not adequately addressed, the resulting treatment policies may fail for certain patient populations. Therefore, to ensure the safe, effective, and equitable deployment of RL systems in real world, rigorous testing, validation, and continuous monitoring are essential.
@@ Appendix A Theoretical Analysis
A.1 The Supplement Analysis of Distribution Shift
Dual Formulation of Model Bias.
The reinforcement learning optimization objective can be equivalently reformulated in its dual form, expressed as $\mathcal{J}_{M}(\pi)=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)]$, where $d^{p,\pi}(s,a)$ denotes the state-action visitation distribution under dynamics $p$ and policy $\pi$. To illustrate the impact of model bias in the dual formulation, we similarly examine the scenario where the training policy $\pi$ is applied within the model dynamics $q$. The training objective $\mathcal{J}(\pi)$ is derived as follows:
$\mathcal{J}(\pi)=\mathbb{E}_{(s,a)\sim d^{q,\pi}}[r(s,a)]=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)\frac{d^{q,\pi}(s,a)}{d^{p,\pi}(s,a)}].$
Model bias can result in inaccuracies in estimating the state-action visitation distribution, subsequently affecting the accuracy of value function estimation. Similarly, if the learned model overestimates or underestimates the state-action visitation distribution, it will lead to corresponding overestimations or underestimations of rewards, thereby distorting the objective.
Trajectory Formulation of Policy Shift.
Similarly, we consider the trajectory formulation of policy shift. We execute the sampling policy $\pi_{c}$ within the environment transition dynamics $p$. Then the training objective $\mathcal{J}(\pi)$ is then given by:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi_{c}}}[R(\tau)]=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\right].$
The policy shift can influence the value estimation of trajectories, imposing implicit constraints on policy learning. For a high-reward trajectory $\tau$, if the current policy $\pi$ closely approximates the optimal policy and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})>\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, it may lead to an underestimation of the return for trajectory $\tau$. Conversely, for a low-reward trajectory $\tau$, if the current policy is relatively effective and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})<\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, this can result in an overestimation of the return for trajectory $\tau$. Both of these impede the effective learning of the policy.
A.2 The Proof of Theorem 1
Theorem A.1.
@@ A.2 The Proof of Theorem 1
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})\text{ , }q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (10)
Then according to (9) and (10) and Jensen’s inequality, we have
$\log\left[\mathcal{J}_{M}(\pi)\right]$ $=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}\right]$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log R(\tau)+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log\frac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\left((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\log r(s_{t},a_{t})+\frac{1}{(1-\gamma)\gamma^{t}}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right)\right]$
$=(1-\gamma)\mathcal{L}(\pi).$
A.3 Classifiers Estimation Derivations

### v3

@@ 5.1. Effectiveness of Shifts-aware Reward
Effectiveness in Complex Tasks. We also conducted experiments on complex tasks. Fig. 2 (c) and (d) shows a visual comparison in the NeoRL task (details in Sec. 5.2.2). SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data by model bias adjustment, whereas the previous state-of-the-art (SOTA) method, MOBILE, imposes weaker constraints. This enables SAMBO to effectively address the distribution shift between offline data and model data during training, thereby allowing for more efficient and reliable model utilization. Benefiting from the policy shift modification, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with both the Offline Data and the Model Data. This enables SAMBO to align with the deployment distribution, effectively addressing the distribution shift between the training data and the data encountered during real deployment, facilitating an efficient connection from training to testing. These ultimately result in an effective resolution to the distribution shift.
Table 2. Normalized average returns in NeoRL benchmark over 4 random seeds. The hightest scores are highlighted in bold.
Task Name BC CQL TD3+BC EDAC MOPO MOBILE SAMBO (Ours)
HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.8 38.4 $\pm$ 3.2
Hopper-L 15.1 16.0 15.8 18.3 6.2 18.2 19.4 $\pm$ 1.3
Walker2d-L 28.5 44.7 43.0 40.2 11.6 23.7 46.1 $\pm$ 5.7
HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 79.0 62.5 $\pm$1.6
Hopper-M 51.3 64.5 70.3 44.9 1.0 43.4 67.6 $\pm$ 12.2
Walker2d-M 48.7 57.3 58.6 57.6 39.9 60.1 63.4 $\pm$ 4.0
HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 71.8 71.6 $\pm$ 13.9
Hopper-H 43.1 76.6 75.3 52.5 11.5 42.3 93.9 $\pm$ 15.2
Walker2d-H 72.6 75.3 69.6 75.5 18.0 71.9 79.2 $\pm$ 0.6
Average 45.4 56.1 54.5 50.7 28.5 51.7 60.2 $\pm$ 6.4
@@
Figure 3. Performance comparison of Hopper tasks in NeoRL benchmark.
5.2. Benchmark Results
5.2.1. D4RL
Datasets. In line with prior research, SAMBO is evaluated using the D4RL benchmark Fu et al. (2020) on the MuJoCo simulator Todorov et al. (2012). The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert). To ensure consistency, we employ the “v2” versions of datasets for standardized evaluation.
Baselines. We evaluate SAMBO against several baseline methods, including model-free methods such as BC Bain and Sammut (1995); Ross et al. (2011) which learns the behavior policy, CQL Kumar et al. (2020) that penalizes Q-values on OOD samples, TD3+BC Fujimoto and Gu (2021) that adopts a BC constraint during optimizing policy, and EDAC An et al. (2021) that penalizes the Q-function based on uncertainty derived from ensemble networks, as well as model-based methods such as MOPO Yu et al. (2020) that penalize rewards based on the uncertainty of the model predictions, COMBO Yu et al. (2021) that extends the penalty of CQL in the model-based setting to regularize OOD samples, TT Janner et al. (2021) that employs a Transformer to model distributions and uses beam search to plan, RAMBO Rigter et al. (2022) that trains the policy and model within a robust framework, and the prior SOTA algorithm MOBILE Sun et al. (2023) that penalizes value function during training by the uncertainty of Bellman Q-function estimates derived from ensemble models.
Comparison Results. Results are presented in Table 1. Each number represents the normalized score, calculated as $100\times$(score - random policy score)/ (expert policy score - random policy score) Fu et al. (2020). SAMBO achieves strong performance on random datasets and demonstrates competitive performance across medium, medium-replay, and medium-expert datasets. Furthermore, SAMBO demonstrates highly competitive average performance, approaching that of the model-free ensemble method EDAC An et al. (2021), and second only to the current SOTA model-based ensemble method MOBILE Sun et al. (2023). However, unlike these methods, SAMBO achieves this competitive performance without relying on value ensemble techniques.
(a) Walker2d-L
(b) Walker2d-M
(c) Walker2d-H
Figure 4. Illustration of ablation experiments in NeoRL benchmark.
5.2.2. NEORL
@@
Datasets. NeoRL Qin et al. (2022) is a benchmark carefully designed to emulate real-world scenarios by generating datasets through a more conservative policy, thereby closely aligning with real-world data-collection scenarios. The datasets, characterized by their narrow scope and limited coverage, present a substantial challenge for offline RL algorithms. In this work, we examine nine datasets across three environments (HalfCheetah-v3, Hopper-v3, Walker2d-v3) and three dataset quality levels (L, M, H), corresponding to low, medium, and high-quality datasets. NeoRL provides varying numbers of trajectories for each task (100, 1000, and 10000), and for consistency, we uniformly selected 1000 trajectories in all our experiments.
@@
Baselines. We compare SAMBO with baseline methods that align with the D4RL experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would be computationally expensive.
@@
Comparison Results. Results are presented in Table 2 and Figure 3. We compare the learning curves of model-based approaches in the NeoRL benchmark. Since MOPO does not have an official implementation on NeoRL, we report its score based on the original NeoRL paper. Compared to other model-based and model-free methods, including ensemble approaches, SAMBO achieves superior performance on the more challenging and realistic NeoRL benchmarks. Moreover, SAMBO consistently outperforms Behavior Cloning (BC), while most baselines show degraded performance compared to BC on certain tasks. In particular, SAMBO demonstrates exceptional performance improvements in the Walker2d and Hopper environments. These results highlight that SAMBO has strong potential for applications on challenging datasets. More experiments and discussions are provided in Appendix D.2.
5.3. Ablation and Tuning Studies
@@
Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Walker2d task in NeoRL benchmark. As shown in Figure. 4, we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), while keeping all other parameters consistent.
The ablations show that, The wo/mb employs policy shift modification to stabilize policy updates. However, it does not address model bias and leads to significant inaccuracies in value estimation, ultimately resulting in poor performance. Conversely, the wo/ps ablation addresses model bias but neglects policy shift adjustments. While this reduces value misestimation, the algorithm remains unstable due to the induced distribution shift between training and deployment. Consequently, wo/ps exhibits lower stability, and its final performance is slightly worse than that of SAMBO. In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating stable updates and addressing value estimation challenges, thereby alleviating suboptimal performance.
Furthermore, we performed parameter tuning experiments on $\alpha$ and $\beta$ to examine the robustness of our method with respect to parameter choices. As illustrated in Figure 5, we fix $\beta=0.01$ in example (a) and $\alpha=0.01$ in example (b). The results indicate that the performance of our algorithm remains generally stable and consistent within a certain range of parameter variations.
(a) Tuning for $\alpha$
(b) Tuning for $\beta$
Figure 5. The tuning experiment was performed on halfchetah-random datasets in D4RL benchmark.
6. Conclusion and Discussion
@@
In this paper, we analyze the distribution shift problem in model-based offline RL, attributing its root causes to model bias and policy shift. Model bias induces a distribution shift among training data originating from different sources, which distorts value estimation, while policy shift results in a distribution shift between training and deployment, hindering policy convergence. We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR). Theoretical and empirical analyses have demonstrated that SAR serves as an efficient surrogate for the true rewards. Building on SAR, we apply it within the model-based offline RL setting, proposing our SAMBO-RL algorithm. Empirical results demonstrate that SAMBO-RL consistently delivers superior or competitive performance, especially on the more challenging NeoRL benchmarks, underscoring SAMBO’s practical effectiveness and validating our theoretical analysis. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for explicit classifier estimation.
@@ Appendix A Theoretical Analysis
A.1. The Proof of Theorem 1
Theorem. 1.
@@ A.1. The Proof of Theorem 1
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})$ (11)
$q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (12)
Then according to (10), (11) and (12) and Jensen’s inequality, we have
$\log[\mathcal{J}_{M}(\pi)]$
$=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}[R(\tau)\tfrac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}]$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[R(\tau)\tfrac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\Big]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\log R(\tau)+\sum_{t=0}^{\infty}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\log\tfrac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})$
$\quad+\sum_{t=0}^{\infty}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\sum_{t=0}^{\infty}((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\sum_{t=0}^{\infty}\gamma^{t}\Big(\log r(s_{t},a_{t})+\tfrac{1}{(1-\gamma)\gamma^{t}}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}$
$\quad+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big)\Big]$
$=(1-\gamma)\mathcal{L}(\pi)$
A.2. Classifiers Estimation Derivations